OpenStudy (anonymous):

Tarzan swings on a 30.0‐m‐long vine initially inclined at an angle of 37^0 with the vertical. What is his speed at the bottom of the swings if he starts from rest?

OpenStudy (anonymous):

I'm stuck, I don't even know where to start

OpenStudy (anonymous):

use conservation of energy.

OpenStudy (anonymous):

So help me solve it?

OpenStudy (anonymous):

first find the height as h = 30 - 30 cos (37) = 6m

OpenStudy (anonymous):

so is that the start?

OpenStudy (anonymous):

then\[\frac{ 1 }{ 2 }mv^{2} = mgh\]\[v^{2}=2gh\]\[v^{2}=2\times9.8\times6\]

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

find v

OpenStudy (anonymous):

so its 10.84 correct?

OpenStudy (anonymous):

whats it measured in?

OpenStudy (anonymous):

m/s